### v1

# Temporal Analysis of Drifting Hashtags in Textual Data Streams: A Graph-Based Application

Abstract
Social media has played an important role since its emergence. People use the internet to express opinions about anything, making social media platforms a social sensor. Initially supported by Twitter, the hashtags are now in use on several social media platforms. Hashtags are helpful to tag, track, and group posts on similar topics. In this paper, we analyze hashtag drifts over time using concepts from graph analysis and textual data streams using the Girvan-Newman method to uncover hashtag communities in annual snapshots. More specifically, we analyzed the #mybodymychoice hashtag between 2018 and 2022. In addition, we offer insights about some hashtags found in the study. Furthermore, our approach can be useful for monitoring changes over time in opinions and sentiment patterns about an entity on social media. Even though the hashtag #mybodymychoice was initially coupled with women’s rights, abortion, and bodily autonomy, we observe that it suffered drifts during the studied period across topics such as drug legalization, vaccination, political protests, war, and civil rights. The year 2021 was the most significant drifting year, in which the communities detected suggest that #mybodymychoice significantly drifted to vaccination and Covid-19-related topics.
keywords:
text drift, hashtag drift, temporal analysis, textual data streams
[…]
[inst1]organization=Programa de Pós-graduação em Informática (PPGIa), Pontifícia Universidade Católica do Paraná (PUCPR),city=Curitiba, state=Paraná, country=Brazil
[…]
[inst2]organization=Instituto Federal de Santa Catarina (IFSC),city=Caçador, state=Santa Catarina, country=Brazil
[…]
[inst3]organization=Universidade Estadual de Ponta Grossa (UEPG),city=Ponta Grossa, state=Paraná, country=Brazil
1 Introduction
The popularization of the Internet made it possible for people to express themselves in several ways via posts on social media using text, audio, videos, and pictures. People can also expose opinions about any topic, e.g., politics, products, and sports, and engage in collective actions and social movements. Social media has also enabled social movements to take place worldwide. For instance, authors in Eltantawy and Wiest (2011) discuss the role of social media in the Egyptian Revolution in 2011. The authors in Harlow (2012) analyzed Guatemalan users’ posts on Facebook groups to evaluate whether social media helped the movement in defense of Rodrigo Rosenberg to go offline in 2010. In addition, the authors in Harlow (2012) named these digital movements ‘Activism 2.0’, also due to Web 2.0, which includes social media networks. Another example of such activism in social media is the ‘Arab Spring,’ also discussed in several papers Smidi and Shahin (2017); Wolfsfeld et al. (2013); Comunello and Anzera (2012).
More recently, during the Covid-19 pandemic, several movements in favor and against vaccination and other aspects of the pandemic, e.g., lockdown, have emerged Sufi et al. (2022). In Menghini et al. (2022), the authors discussed the use of #mybodymychoice, a feminist slogan to defend women’s rights, used for opposing Covid-19 vaccination campaigns.
Twitter is one of the most popular social media platforms. Around 500 million tweets are posted per day in 2022111https://www.dsayce.com/social-media/tweets-day/. Each post is known as tweet. Its ease of posting and sharing information made it popular also for famous people, such as artists, athletes, and politicians. As of 2017, Twitter limits the post to 280 characters in its free version. In addition, people can tag their posts by using hashtags, leading to a post categorization. Researchers that need to evaluate information from social sensors, e.g., public opinion, resort to data from social media Suprem et al. (2019).
Inspired by the authors in Menghini et al. (2022), we also analyzed the Twitter stream regarding the hashtag #mybodymychoice, comprising the years between 2018 and 2022. The authors in Menghini et al. (2022) demonstrated how the hashtag mentioned before, originally used to defend the right to abortion and other bodily autonomy aspects, drifted over time, being also used in different topics. Our work technically analyzes the context evolution of the hashtag, originating from a textual data stream corresponding to #mybodymychoice without judging the validity of the related social causes or the drifts of this hashtag. In addition, different from Menghini et al. (2022), which split the stream using two events known beforehand, we leveraged the Girvan-Newman community detection algorithm Girvan and Newman (2002) between years to evaluate the communities found.
This paper aims to analyze the drifts in the hashtag stream of #mybodymychoice using time slices corresponding to years between 2018 and 2022, inclusive. We applied the Girvan-Newman algorithm for community detection after each year was finished. We analyzed the most relevant, i.e., the biggest, communities by evaluating the hashtags and their contexts. We are interested in analyzing the #mybodymychoice’s distancing from its original use or context. In addition, we listed the five most frequent hashtags per year. We consider drifts to have occurred whenever the hashtag is used in different contexts than its original. Therefore, our research questions are (RQ1) “How the hashtag #mybodymychoice evolved from 2018 to 2022?”, and (RQ2) “What were the most significant hashtag drifts over time, considering the graph communities detected?”. RQ1 aims to uncover drifts and their importance regarding a conceptual similarity to the original use of #mybodymychoice. RQ2 aims to find which topics other than women’s rights, abortion, and bodily autonomy were frequent in the period analyzed.
The contribution of this work is two-fold: (i) the analysis of the hashtag stream respecting the textual data stream concepts and using graph-based approaches, including community detection between years, and (ii) to the best of our knowledge, this is the first time the Girvan-Newman algorithm was applied to a hashtag graph generated in a streaming fashion. A by-product of this work is the dataset containing hashtag stream regarding the #mybodymychoice hashtag between 2018 and 2022. This work may be relevant to applications that monitor sentiments or opinions regarding an entity on social media to recognize changes in patterns and behaviors.
This work is organized as follows. Section 2 presents the context background, including hashtags, textual data stream mining, graphs, and community detection. Section 3 describes the methodology used in the analysis, including the dataset collection and hashtags treatment. Section 4 presents the graph generation and implementation. Section 5 analyzes the hashtags that co-occur with #mybodymychoice to infer topics using community detection in graphs, considering the years between 2018 and 2022. Finally, Section 6 concludes this work and provides future directions.
2 Background
This section introduces the concepts to understand the approach. We present hashtags and their uses, textual data stream mining concepts, graphs, and community detection in graphs.
2.1 Hashtags
Hashtags are words starting with the hash character (#) used for social media post categorization. Hashtags, under the name of channel tags, were elaborated by Chris Messina in 2007 with the idea of creating groups on Twitter [^2] Available at: https://factoryjoe.com/2007/08/25/groups-for-twitter-or-a-proposal-for-twitter-tag-channels/, accessed on June 20th, 2023. Twitter adopted hashtags as a feature in 2009 Scott (2015), which proved helpful and, later, was also integrated into other social media platforms, such as Facebook [^3] Available at: https://about.fb.com/news/2013/06/public-conversations-on-facebook/. Accessed on June 20th, 2023. in 2013.
Twitter post categorization or grouping leads to virtual community creation. This enables discussions on several aspects, and sometimes, online social movements go offline Smidi and Shahin (2017). An example of a hashtag used in social discussion is #mybodymychoice. As pointed out in Menghini et al. (2022), the hashtag #mybodymychoice initially regarded women’s right to abortion and bodily autonomy. However, the authors in Menghini et al. (2022) found that between 2018 and 2021, this hashtag drifted to other topics, such as politics and civil rights.
This paper expands the analysis performed in Menghini et al. (2022) to 2022, considering hashtags co-occurrence and community detection via temporal graphs and textual streaming analysis. In our analysis, we intend to trace back the events that used particular hashtags to understand the hashtags’ occurrences. Therefore, this section presents textual data stream mining and its characteristics, graphs and their notation, and techniques for community detection in graphs.
2.2 Textual Data Stream Mining
Data streams are “an algorithmic abstraction that allows for real-time analytics” Gama et al. (2014). Data arrive sequentially (one by one or in small batches) and quickly in streaming settings. In addition, a data stream may be infinite and ephemeral, meaning that the data distribution may change over time, thus giving rise to a phenomenon named concept drift Bifet et al. (2018); Gama et al. (2014).
Therefore, to learn from data streams, i.e., data stream mining, some capabilities desired for machine learning methods include Bifet et al. (2018); Gama et al. (2014): (a) learning quickly from the data as it arrives, (b) discarding data within a short time after learning from it, (c) performing single-pass processes, and (d) detecting and adapting to concept drifts.
Textual data streams are specializations of traditional data streams Thuma et al. (2023), i.e., it comprises a sequence of texts arriving continuously at a fast pace. However, processing texts in textual data stream scenarios is more challenging given the natural language processing (NLP)-related activities regarding, for example, vocabulary maintenance, token standardization, and processing texts only once.
Our proposed approach focuses on processing textual data streams while accounting for the above-mentioned constraints. The exception regards graph community detection, detailed in Section 2.4, which is performed on a year-basis, and, thus, should be considered an offline step.
2.3 Graph
A graph $G$ is a collection of edges $E$ and vertices $V$. Each vertex can be linked to another vertex by an edge. Generally, vertices are visually represented as circles, and the edges as lines between vertices, as depicted in the example graph in Fig. 1. Edges are directed if the direction between vertices matters and are undirected otherwise. Edges may also be coupled with weights to denote additional information about the relationship between connected vertices.
Figure 1: An example graph with four vertices (#hashtags) and three undirected edges (lines).
Graphs are relevant to denote entities and the relationships between them, being very useful in text analysis. For instance, to find relevant words in sentences, the vertices are the words, and the edges are their relationships Castillo et al. (2017). In our proposal, vertices denote hashtags, and edges denote the co-occurrence between the linked hashtags. The edges are not weighted.
2.4 Community Detection in Graphs
Community detection in graphs groups nodes (vertices) with frequent links into a community. Criminal community detection Sarvari et al. (2014), cattle trade movement Cardoso et al. (2018), and disease spread dynamics Salathé and Jones (2010) are applications that leverage community detection. Several community detection methods are available in the literature, such as the Louvain method Blondel et al. (2008), the Leiden algorithm Traag et al. (2019), and the Girvan-Newman algorithm Girvan and Newman (2002).
Similar to clustering, community detection demands the graph partition into densely connected nodes and nodes from a certain community to be far from those of other communities Blondel et al. (2008). According to Traag et al. (2019), the Louvain method obtains weakly connected communities. To handle this characteristic, the authors in Traag et al. (2019) added one more phase to the Louvain algorithm to obtain better-connected communities.
The Girvan-Newman algorithm Girvan and Newman (2002) for community detection on graphs performs an iterative edge elimination. To consider an edge for elimination, it must be part of the highest number of shortest paths between nodes, i.e., betweenness centrality. After removing these edges, the communities are obtained. More specifically, the algorithm works in four steps Girvan and Newman (2002): (a) for all edges in the graph, compute the betweenness; (b) remove the edge with the highest betweenness; (c) for all edges affected by the edge removal in step (b), re-compute the betweenness; (d) return to step (b) until there is no edge remaining. We resorted to the Girvan-Newman algorithm because it generates consistent communities Sheng et al. (2020). However, its trade-off lies in its complexity, which makes it not scalable for large networks Sheng et al. (2020).
Considering that we are using a textual data stream limited by a window and that the communities are detected in an offline step, community quality is the most important. Therefore, the Girvan-Newman algorithm’s complexity is not a concern in this particular scenario.
3 Methodology
In this section, the data collection is detailed, discussed, and the obtained dataset is described. In addition, we present the hashtags treatment.
3.1 Dataset description
The dataset regarding the #mybodymychoice hashtag comprises tweets from February 11th, 2018, to December 31st, 2022. Data was acquired using the Twitter API [^4] https://developer.twitter.com/en/docs/twitter-api with the “(#$(\#mybodymychoice)\ -is:retweet$” query, which ignores retweets (re-posts). In addition, we performed the collection respecting Twitter’s policies for that. We ignored retweets since they could bias our analysis by inflating the use of particular hashtags. The dataset totaled 255,131 valid tweets. Figure 2 depicts the tweet distribution over the analyzed period.
Figure 2: Number of tweets distributed across the period evaluated.
The peaks in Figure 2 provide interesting insights. To provide information to investigate the peaks, we collected the five most frequent hashtags in the months of interest, i.e., May 2019, August and September 2021, and May 2022. In the first peak, i.e., May 2019, the five most frequent hashtags, together with their respective frequencies, were #prochoice (1090), #womensrightarehumansrights (863), #womensrights (728), #abortionisawomansright (713), and #stopthebans (552). Investigating this peak, we retrieved that in May 2019, several protests against abortion-related restrictions happened in the United States [^5] Available at: https://www.npr.org/2019/05/21/725410050/across-the-country-protesters-rally-to-stop-states-abortion-bans. Accessed on May 30th, 2023..
Repeating the analysis for the second peak in August 2021, the five most frequent hashtags and respective frequencies were: #novaccinepassports (519), #freedom (443), #novaccinepassportsanywhere (397), #novaccinepassport (382), and #covid19 (330). We can see that this peak starts earlier than August. It matches the release of the European Union Digital Covid Certificate, released on July 1st, 2021666Available at: https://commission.europa.eu/strategy-and-policy/coronavirus-response/safe-covid-19-vaccines-europeans/eu-digital-covid-certificate_en. Accessed on May 30th, 2023.. In September, the most frequent hashtags were #womenrights (1321), #texaswaronwomen (1282), #abortionishealthcare (1268), #prochoice (1231), and #novaccinepassports (1187). The second hashtag provides interesting information about the “Texas war on women”. In Texas, protests against the state abortion-related laws emerged in this period [^7] Available at: https://www.theguardian.com/global-development/2021/sep/18/texas-anti-abortion-law-shows-terrifying-fragility-of-womens-rights-say-activists. Accessed on May 30th, 2023..
The third peak, between May and August 2022, reflects an important event regarding the #mybodymychoice. In June 2022, the US Supreme Court invalidated the Roe vs. Wade decision (which favored abortion in 1973) [^8] Available at: https://www.supremecourt.gov/opinions/21pdf/19-1392_6j37.pdf. Accessed on May 30th, 2023..
In summary, the data collected provides several insights and reflects the past regarding the hashtag #mybodymychoice. The link for downloading the dataset and source code will be made upon paper acceptance.
3.2 Hashtags treatment
[…]
4 Experimental Protocol
This section presents the graph generation process and its implementation, including the pseudocodes.
4.1 Graph generation
We generated graphs using their intrinsic characteristics: each node corresponded to a hashtag, and each edge between two nodes corresponded to their co-occurrence. We considered a hashtag valid once it appears in the dataset at least five times to ease the analyses and avoid noise. Considering the textual data stream setting, we perform the computation on-the-fly. It means that the graph generation process respects the textual data stream constraints. The only exception is the community detection performed at the end of each year (regarding the timestamp of the textual data stream).
In addition, to limit our analysis to the most recent and important hashtags, we used a 200-size window and a hashtag aging mechanism. In our approach, each hashtag has an age attribute. This attribute controls the hashtags to be considered in the analysis. Whenever a new hashtag is to be inserted, if the window is full, the oldest hashtag in the window is removed to accommodate the new hashtag. If the window is not full, the hashtag is added. In each iteration, the hashtags’ age is increased, except for the recently received hashtag.
4.2 Implementation
We implemented the code in Python 3.9. We plotted the graphs with the help of the Pyvis [^9] https://pyvis.readthedocs.io/en/latest/ and NetworkX libraries [^10] https://networkx.org/. The link for downloading the source code and dataset will be available upon paper acceptance. Algorithm 1 depicts the pseudocode. The algorithm receives a textual data stream $TS$ and creates a new Graph with a window size of 200 and a minimum frequency of five. Line 2 starts the stream iteration; line 3 obtains the next text (tweet); line 4 extracts a list of hashtags, which is sent to the add_node graph’s method, detailed in Algorithm 2.
Algorithm 1 Textual Data Stream Processing
1:Textual data stream $TS$
2:$g\leftarrow new\ Graph(200,5)$
3:while $TS\neq\O$ do
4:     $text\leftarrow TS.get\_next\_text()$
5:     $hashtags\leftarrow text.get\_hashtags()$
6:     $g.add\_node(hashtags)$triangleright Algorithm 2
7:end while
The Algorithm 2 details the add_node method. In line 1, all the hashtags’ age is incremented. In line 3, for each hashtag in the received list, we check if the hashtag is not in the vocabulary (vocab), i.e., it is not in the graph, if it is in the pregraph, and if its frequency reached five simultaneously. Pregraph is a graph’s internal structure that stores the hashtags until they reach the minimum defined frequency. Also, g.vocab is the structure in which we control the existence of nodes, i.e., hashtags. Once a given hashtag reaches the minimum frequency, it can be added to the graph in line 4. In line 5, the hashtag is removed from the pregraph. If the condition in line 3 is not met, it is checked if the hashtag is in the pregraph. If the condition is met, its frequency is incremented. Next, we evaluate the remaining hashtags as connections. In line 10, we call the variable other_hashtags. The same process of checking existence in g.vocab and pregraph is performed between lines 12 and 18. In line 19, we connect the hashtag and the other_hashtag. In lines 20 and 21, we reset the age of hashtag and other_hashtag. The age reset is important so the window can identify and remove stale hashtags when opportune.
Algorithm 2 Add node mechanism
1:Hashtags list $hashtags$
2:Graph $g$
3:$g.grow\_old()$
4:for each $hashtag$ in $hashtags$ do
5:     if $hashtag\notin g.vocab$ and $hashtag\in g.pregraph$ and $g.pregraph(hashtag)\geq 5$ then
6:         $g.create\_node(hashtag)$
7:         $g.pregraph.remove(hashtag)$
8:     else if $hashtag\in g.pregraph$ then
9:         $g.pregraph(hashtag).increment()$
10:         continue
11:     end if
12:     $other\_hashtags\leftarrow hashtags-hashtag$
13:     for each $co\_hashtag$ in $other\_hashtags$ do
14:         if $co\_hashtag\notin g.vocab$ and $co\_hashtag\in g.pregraph$ and $g.pregraph(co\_hashtag)\geq 5$ then
15:              $g.create\_node(co\_hashtag)$
16:              $g.pregraph.remove(co\_hashtag)$
17:         else if $co\_hashtag\in g.pregraph$ then
18:              $g.pregraph(co\_hashtag).increment()$
19:              continue
20:         end if
21:         $hashtag.add\_connection(other\_hashtag)$
22:         $hashtag.age\leftarrow 0$
23:         $other\_hashtag.age\leftarrow 0$
24:     end for
25:end for
It is important to mention that we do not include the hashtag #mybodymychoice because all the nodes would be connected to the #mybodymychoice’s node.
5 Results
This section shows the results obtained sliced by year. As mentioned, we considered the most relevant communities for analysis. It is crucial to highlight that the graphs are not reset between years: they are constructed and updated incrementally. However, the colors have no correspondence across the years. For example, the cluster in green in 2018 may not be the same in green in 2019.
5.1 Analysis of 2018
Figure 3 depicts the remaining hashtags after the process described in Section 4. Each color represents a community detected by the Girvan-Newman method.
Figure 3: Relationship between hashtags, considering their co-occurrences. The colors represent the communities generated by the Girvan-Newman method. Best viewed in color.
The largest community, i.e., in green, regards hashtags such as #safeabortion, #healthcare, #reprorights, #birthcontrol, and #plannedparenthood. Therefore, this community is closely related to its original context: feminism and bodily autonomy, also considering abortion aspects. Fig. 4 partially shows the green community zoomed in. The community in purple regards different aspects, such as #feminism, #vibes, #tattoo, and #gaypower. The terms are mixed, and it is impossible to infer an exact context. Fig. 5 partially depicts this community.
Figure 4: Part of the green community from 2018.
Figure 5: Part of the purple community from 2018.
Conversely, the community in orange, i.e., the community at the bottom left, considers hashtags that refer to civil rights, for example, #blacklivesmatter, #noracistwall, #resistance, and #votethemout. The community in yellow regards hashtags such as #youdontownme, #fridaythoughts, and #impeachtrump. Although they are closely connected, inferring a context is impossible. Figs. 6 and 7 show the communities in orange and yellow, respectively. Some labels are not shown due to the existence of swear words.
Figure 6: Part of the orange community from 2018.
Figure 7: Part of the yellow community from 2018.
The community in blue considers the use of drugs. The hashtags that prevail are: #novictimnocrime, #recreationalcannabis, #legalisecannabis, and #cannabisismedicine. Fig. 8 depicts the blue community.
Figure 8: Blue community from 2018.
In summary, regarding 2018, the hashtag #mybodymychoice was used in different contexts than its original one, drifting to civil rights, the use of drugs, and politics. However, the largest community regards the original context: feminism, abortion, and women’s bodily autonomy. The five most frequent hashtags in 2018 and respective frequencies were: #prochoice (987), #repealthe8th (532), #womensrights (482), #metoo (379), and #fbr (339).
5.2 Analysis of 2019
Fig. 9 shows the communities detected by the Girvan-Newman method from the graph updated until the end of 2019. The biggest community is the purple one, partially depicted in Fig. 10. This community includes hashtags across different topics such as #humanrights, #vaccination, #forcedvaccination, #drugs, #prochoice, #climateemergency, #greennewdeal, and #presidentberniesanders. Therefore, at first sight, it ranges from human rights to politics. According to his official website [^11] Available at: https://berniesanders.com/issues/reproductive-justice-all/. Accessed on May 27th, 2023, Bernie Sanders considers abortion a constitutional right. Therefore, it makes sense that the hashtag #presidentberniesanders and other hashtags which seem to be related to Bernie Sanders’ campaign for president in the United States appear together with #prochoice and #humanrights. In addition, #greennewdeal is a topic of his campaign related to environmental issues. In essence, this community regards the civil rights and campaign topics of Bernie Sanders.
Figure 9: Relationship between hashtags, considering their co-occurrences regarding 2019. The colors represent the communities generated by the Girvan-Newman method. Best viewed in color.
Figure 10: Part of the purple community from 2019.
The community in blue also stands out. It contains hashtags such as #mykidsmychoice, #votethemout, #firstamendment, #votenos2173, and #killbills2173. The hashtag #votethemout is probably related to the unacceptance of the government in charge in 2019. The hashtag #firstamendment prevents the US government from making laws prohibiting religions and their exercise [^12] Available at: https://www.whitehouse.gov/about-the-white-house/our-government/the-constitution. Accessed on May 27th, 2023.. In addition, #killbills2173 and #votenos2173 regard Bill S2173131313Available at: https://www.billtrack50.com/BillDetail/966141. Accessed on May 31st, 2023., which intended to eliminate the religious exemption for mandatory childhood vaccination. Therefore, this community combines hashtags related to religion, politics, and vaccination. Fig. 11 shows the major part of the community in blue.
Figure 11: Part of the blue community from 2019.
The community in blue-gray, i.e., at the center top, considers hashtags such as #prolife, #adoption, #endabortion, #rescuethepreborn, and #restorefamilies. This community is undoubtedly anti-abortion. Conversely, the dark-orange-colored community contains hashtags such as #nojustice4men, #womenempowerment, #nolaw4man, #norepublic4men, and #genderbiasedlaws. It seems that this community regards feminism or might be considered hate speech. Figs. 12 and 13 shows the blue-gray and the dark-orange communities, respectively.
Figure 12: Part of the blue-gray community from 2019.
Figure 13: Part of the dark-orange community from 2019.
It is also possible to notice that new communities are growing in 2019 but still have a few components. In addition, we see that the #mybodymychoice regards vaccination, feminism, civil rights, and politics in this time slice. The five most frequent hashtags in 2019 and respective frequencies were: #prochoice (2451), #womensrights (1384), #womensrightsarehumanrights (1207), #abortionisawomansright (916), and #abortion (905).
5.3 Analysis of 2020
Fig. 14 depicts the communities discovered by the Girvan-Newman algorithm considering the hashtags until the end of 2020. Two big communities draw attention: those in blue and yellow.
Figure 14: Relationship between hashtags, considering their co-occurrences regarding 2020. The colors represent the communities generated by the Girvan-Newman method. Best viewed in color.
Exploring the biggest community, i.e., the blue-colored community, we see hashtags such as #abortion, #abortionrights, #reproductivehealth, #womensrights, #abortionishealthcare, #abortionwithoutborders, and #abortionisnotacrime. In addition, hashtags in languages other than English expressing the same stance are present, i.e., #abortolegal (“legal abortion” in Portuguese or Spanish), #avortement (“abortion” in French), and #abortolegalya (“legal abortion now” in Spanish). However, since #abortion is present in this community, opinions against abortion are present, for example, the hashtags #rescuethepreborn, #chooselife, #savethebabyhumans, and #endabortions. Therefore, this community, depicted in Fig. 15, combines hashtags in favor and against abortion.
Figure 15: Part of the blue community from 2020.
In the yellow community, the terms draw attention. In 2019, the first terms related to regular vaccination appeared. However, in the 2020 analysis, several terms regarding Covid-19 vaccination emerged.
In November 2019, a new pneumonia-related disease appeared in China, later named Covid-19, after coronavirus disease and the year this disease emerged. Due to its ease to spread and severity, the World Health Organization (WHO) declared Covid-19 a global pandemic in March 2020 Cucinotta and Vanelli (2020). Therefore, WHO suggested measures to prevent the Covid-19 spread, such as wearing masks, keeping distance from other people, and washing hands [^14] Available at: https://www.who.int/westernpacific/emergencies/covid-19/information/transmission-protective-measures. Accessed on May 30th, 2023.. The US Food and Drugs Association (FDA) authorized the first Covid-19 vaccine for emergency use in December 2020151515Available at: https://www.fda.gov/news-events/press-announcements/fda-takes-key-action-fight-against-covid-19-issuing-emergency-use-authorization-first-covid-19. Accessed on May 30th, 2023..
Thus, 2020 was a year of changes in daily life, and the hashtags demonstrate that. Hashtags such as #nomoremasks, #masksdontwork, #lockdownsdontwork, #standup, #medicalfreedom, #covidiot, #pharma, and #covid19 appeared, demonstrating opinions against the masks, vaccines, and lockdowns. Therefore, the yellow community, depicted in Fig. 16, is majorly related to the Covid-19 pandemic.
Figure 16: Part of the yellow community from 2020.
The community in dark orange is also related to the Covid-19 pandemic, most specifically against governmental measures and governments themselves. Some of the hashtags are #scamdemic (combination of scam and pandemic), #thegovernmentisthevirus, #iwillnogetvaccinated, #iwillnocomply, #scamdemic, and #wehavebeenliedto. Fig. 17 depicts this community. We suppressed a few hashtags due to the existence of swearwords.
Figure 17: Part of the dark orange community from 2020.
Finally, the community in green regards circumcision, which regards the foreskin removal through surgery. Some of the hashtags are #myforeskinmychoice, #circumcision, #saynotocircumcision, #bornthisway, and #naturalbaby. The hashtags demonstrate opinions against circumcision. We can also see new hashtags without communities appearing. Fig. 18 depicts this community.
Figure 18: The green community from 2020.
In summary, although there are hashtags related to the original aim of #mybodymychoice, it is possible to see that the hashtag #mybodymychoice shifted to topics other than abortion and women’s bodily autonomy. Covid-19-related hashtags emerged in 2020. The five most frequent hashtags in 2020 and respective frequencies were: #prochoice (134), #abortion (64), #strajkkobiet (54), #votenos2173 (53), and #prolife (50).
5.4 Analysis of 2021
The hashtags relationships collected at the end of 2021 are depicted in Figure 19. The community in yellow contains hashtags such as #dictator, #eucovidcertificate, #antivax, #novaccinepassportanywhere, #fighttyrany, #freedomofchoice, #novax, and #covidscam. We see that most hashtags are against vaccination, vaccine passport, and in favor of having the choice to have the vaccine or not. Although these hashtags essentially express opinions against vaccination or its governmental control, some hashtags in favor of vaccination also appear. For example, #getvaccinated and #getvaccinatednow appear among the contrary hashtags. Considering the analysis of the studied period, i.e., between 2018 and 2022, this is the first time the biggest community is not related to abortion and women’s bodily autonomy. Fig. 20 shows part of the yellow community.
Figure 19: Relationship between hashtags, considering their co-occurrences regarding 2021. The colors represent the communities generated by the Girvan-Newman method. Best viewed in color.
Figure 20: Part of the yellow community from 2021.
The blue community considers hashtags from different topics but majorly regards abortion and women’s bodily autonomy. Some of the hashtags are: #abortionishealthcare, #womensrights, #plannedparenthood, #proabortion, and #reproductiverights. In addition, hashtags such as #roevwade and #roe appear. These hashtags regard a decision taken by the US Supreme Court in favor of women choosing to have an abortion, which previously was criminalized [^16] Available at: https://supreme.justia.com/cases/federal/us/410/113/. Accessed on May 30th, 2023.. Unrelated hashtags, such as #animalrights, #storytelling, #writers, and #wordpress also appear. Fig. 21 partially depicts the blue community.
Figure 21: Part of the blue community from 2021.
The community in dark orange demonstrates opinions against the Covid-19 vaccine, specifically Sinovac, with the hashtags #sinopharm, #hypocrite, #corrupt, and #sinovac. The community in blue-gray considers opinions against lockdowns and the Canadian Prime Minister in charge, Justin Trudeau. The hashtags in this community are #fightforfreedom, #trudeauworstpmever, #nomorelockdowns, and #crimesagainsthumanity. Both communities are depicted, respectively, in Figs. 23 and 23.
Figure 22: Part of the dark-orange community from 2021.
Figure 23: Part of the blue-gray community from 2021.
In summary, the analysis of 2021 highlights the first time in the analyzed period that the biggest detected community does not relate to abortion and women’s bodily autonomy. It shows that, at least temporarily, the hashtag #mybodymychoice drifted to vaccination and Covid-19-related topics. The five most frequent hashtags in 2021 and respective frequencies were: #novaccinepassports (3694), #prochoice (2736), #novaccinepassportsanywhere (2605), #freedom (2524), and #novaccinemandates (2305).
5.5 Analysis of 2022
Figure 24 depicts the communities detected by the Girvan-Newman method at the end of 2022. Compared to the previously analyzed years, more communities were detected.
Figure 24: Relationship between hashtags, considering their co-occurrences regarding 2022. The colors represent the communities generated by the Girvan-Newman method. Best viewed in color.
Starting with the community in blue, hashtags related to abortion and women’s bodily autonomy appear. For example, #plannedparenthood, #abortionpills, #endgunviolence, #feminism, #abortionishealthcare, #abortionrightsarehumanrights, #myuterusmychoice, #prochoice, #prolife, and #abortion. However, hashtags unrelated to the abortion topic also emerge, such as #blessed, #covid, #endgunviolence, and #glorytoukraine. The latter two hashtags possibly relate to the Russian-Ukrainian conflict, which started in November 2022. Fig. 25 partially shows the blue community.
Figure 25: Part of the blue community from 2022.
The community in yellow contains hashtags related to Covid-19. For example, #novaccinemandates, #vaccineinjuries, #vaccinesideeffects, #informedconsent, #standupforfreedom, #vaccinedisaster, #lockdowns, and #bigpharma appear in this community. Different from 2021, in 2022, hashtags related to vaccine side effects and injuries appear more frequently than hashtags against vaccination in a Covid-19-related group. The dark-green community also considers Covid-19-related hashtags such as #greenpass, #nojabnojob, #italianblackmail, #novax, and #pfizergate. It is closely related to Covid-19 vaccination since Green Pass is the name on which the EU Digital Covid Certificate is known in Italy [^17] Available at: https://italygreenpass.com/. Accessed on May 30th, 2023.. A dark-orange community also emerged containing hashtags related to Covid-19, such as #yourbodyyourchoice, #thejab, and #covidisntover. These communities are partially depicted in Figs. 26, 27, and 28 respectively.
Figure 26: Part of the yellow community from 2022.
Figure 27: Part of the dark-green community from 2022.
Figure 28: Part of the dark-orange community from 2022.
The community in blue-gray regards sexual topics, such as #nudes, which consider nudity. Also, the community in dark red is similar to the dark-orange community, having hashtags such as #braless and #freethenipples. The community in light orange combines different hashtags, e.g., #mutilation, #genital, #cutting, #children, and #circumcision. This last community is related to the circumcision topic, probably against it, since other hashtags, such as #bornperfect, #trauma, and #ididnotconsent are also present, as can be seen in Fig. 29.
Figure 29: Light orange community from 2022.
In summary, the biggest communities in the analysis of 2022 regard mostly Covid-19, abortion, and women’s bodily autonomy-related topics. Other topics, such as politics and sexual-related, also appear. We also highlight that hashtags related to the Russian-Ukrainian conflict and vaccination side effects emerged. The five most frequent hashtags in 2022 and respective frequencies were: #roevswade (17719), #roevwade (11179), #womensrights (8750), #prochoice (7987), and #womensrightsarehumanrights (7961).
6 Conclusion and Future Works
In this paper, we analyzed the hashtag #mybodymychoice between 2018 and 2022. We generated graphs based on hashtags and their co-occurrences over time and performed annual analyses. The generated graphs went through the Girvan-Newman method for community detection, and the hashtags were discussed to infer topics to the communities and verify whether the hashtag #mybodymychoice drifted.
We verified that the #mybodymychoice hashtag was used in topics related to its original use through time: abortion and women’s bodily autonomy. However, considering the analyzed period, unrelated topics used the aforementioned hashtag. For instance, topics such as politics, drugs, Covid-19, and related topics, e.g., vaccination and vaccine passport, used this hashtag, mostly in 2021, where the biggest community found was related to Covid-19.
More specifically, we aimed to address the research questions presented in Section 1. Considering (RQ1) “How the hashtag #mybodymychoice evolved from 2018 to 2022?”, we concluded that the #mybodymychoice in the period analyzed always showed uses other than its original context, i.e., women’s rights, abortion, and bodily autonomy. In 2018, the largest community regarded the hashtag’s original context. In 2019, the largest community included hashtags related to vaccination and political campaigns. In 2020, the biggest community regarded the hashtag’s original context, using hashtags in languages other than English. Regarding 2021, the yellow community draws attention since it relates to Covid-19, not the original context. In 2022, the biggest community again regards women’s rights, abortion, and bodily autonomy. New hashtags related to the Russian-Ukrainian conflict also appear.
Regarding (RQ2) “What were the most significant hashtag drifts over time, considering the graph communities detected?”, we concluded that 2021 provided the most significant drift, despite the constant minor drifts of the hashtag #mybodymychoice. The hashtag drift in 2021 regarded a crucial moment of humanity: the Covid-19 pandemic. Thus, several governmental decisions were questioned by a fraction of society, such as the mandatory vaccination and the use of a document to prove the vaccination, known as the vaccine passport.
As future works, we intend to update and evaluate the described dataset using incremental algorithms for community detection. In addition, algorithms for topic inference from the hashtags may be used.
CRediT authorship contribution statement
Cristiano M. Garcia: Conceptualization, Methodology, Software, Validation, Investigation, Data Curation, Writing - Original Draft, Presentation. Alceu de S. Britto Jr.: Methodology, Validation, Writing - Review & Editing, Supervision. Jean P. Barddal: Methodology, Validation, Writing - Review & Editing, Supervision.
Data Availability
Under Twitter’s policies, the data are publicly available at: https://github.com/cristianomg10/temporal-analysis-of-drifting-hashtags-in-textual-data-streams-a-graph-based-application.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.